In this talk we discuss a construction of indecomposable p-algebras with exponent
  p and large index. This is closely connected with the study of degeneracy in
 matrices defining abelian crossed products and torsion in CH^2 of Severi-Brauer
    varieties as developed by Karpenko and others. These connections are not
        completely understood and some open questions will be discussed.